### v2

# New AGN diagnostic diagrams based on the [OIII] auroral line

New AGN diagnostic diagrams based on the [OIII] auroral line
The James Webb Space Telescope (JWST) is revolutionizing our understanding of black hole formation and growth in the early Universe. However, JWST has also revealed that some of the classical diagnostics, such as the BPT diagrams and X-ray emission, often fail to identify active galactic nuclei (AGN) at high redshift and low metallicity. Here we present three new rest-frame optical diagnostic diagrams to identify narrow-line Type II AGN, leveraging the [OIII]auroral line, which has been detected in several JWST spectra. Specifically, we show that high values of the [OIII]Hratio provide a sufficient (but not necessary) condition to identify the presence of an AGN, both based on empirical calibrations (using local and high-redshift sources) and a broad range of photoionization models. These diagnostics are able to separate much of the AGN population from Star Forming Galaxies (SFGs). This is because the average energy of AGNās ionizing photons is higher than that of young stars in SFGs, hence AGN can more efficiently heat the gas, therefore boosting the [OIII]line. We also found independent indications of AGN activity in some high-redshift sources (zĀæ4) that were not previously identified as AGN with the traditional diagnostics diagrams, but that are placed in the AGN region of the diagnostic presented in this work. We note that, conversely, low values of [OIII]Hcan be associated either with SFGs or AGN excitation. We note that the fact that strong auroral lines are often associated with AGN does not imply that they cannot be used for direct metallicity measurements (provided that proper ionization corrections are applied), but it does affect the calibration of strong line metallicity diagnostics.
Key Words.:
galaxies: active ā galaxies: high-redshift - galaxies: ISM1 Introduction
Thanks to the successful launch of the James Webb Space Telescope (JWST; Gardner etĀ al., 2023; Rigby etĀ al., 2023), we can now investigate with high resolution and unprecedented sensitivity both the photometric and spectroscopic properties of galaxies up to (Curtis-Lake etĀ al., 2022; Robertson etĀ al., 2022). Within this context, recent studies, exploiting different kinds of JWST data, have revealed a large population of AGN at high redshift (Kocevski etĀ al., 2023; Ćbler etĀ al., 2023; Matthee etĀ al., 2023; Maiolino etĀ al., 2024b, 2023a; Greene etĀ al., 2023; BogdĆ”n etĀ al., 2023; Goulding etĀ al., 2023; Kokorev etĀ al., 2023; Furtak etĀ al., 2023; Juodžbalis etĀ al., 2024b; Scholtz etĀ al., 2023a; Chisholm etĀ al., 2024), suggesting a higher AGN fraction at than previously expected. For instance, Yang etĀ al. (2023), taking advantage of the JWST-MIRI photometry of the Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein etĀ al., 2022), investigated the AGN population using spectral energy distribution (SED) modelling, and found a black-hole accretion rate density (BHARD) at dex higher than what was expected from previous X-ray studies (Vito etĀ al., 2016, 2018). Also Maiolino etĀ al. (2023a) and Harikane etĀ al. (2023), by selecting broad-line AGN (BLAGN, or TypeI) among NIRSpec spectra (Ferruit etĀ al., 2022; Jakobsen etĀ al., 2022), have found significant AGN excess at with respect to the AGN luminosity functions derived using X-ray data (Giallongo etĀ al., 2019).
The X-ray emission, produced in AGN by inverse Compton scattering of the UV photons coming from the SMBH accretion disc, has been traditionally used to select AGN. Most of the high-z AGN selected using JWST spectroscopy lie in fields covered by some of the deepest X-ray observations ever performed, like the 7Ms Chandra Deep Field South (CDFS) (Luo etĀ al., 2017), Chandra Deep Field North (CDFN) (Xue etĀ al., 2016) and AEGIS-XD (Nandra etĀ al., 2015). However, most of these newly discovered high-z AGN lack any X-ray emission (e.g. Mazzolari et al., in prep., Scholtz etĀ al., 2023a; Ćbler etĀ al., 2023; Maiolino etĀ al., 2024a; Kocevski etĀ al., 2023; Wang etĀ al., 2024; Juodžbalis etĀ al., 2024a), with only very few exceptions (Goulding etĀ al., 2023; Kovacs etĀ al., 2024; Kocevski etĀ al., 2023). This makes the picture even more difficult to interpret and suggests different scenarios such as Compton thick obscuration (Gilli etĀ al., 2022) or intrinsic X-ray weakness (Simmonds etĀ al., 2016).
The difficulties in selecting high-z AGN are even more important for TypeII or narrow-line AGN (NLAGN). Different observational works have already demonstrated that some of the traditional well-known AGN emission-line diagnostic diagrams, like the BPT (Baldwin etĀ al., 1981) diagram, are no longer effective in the high-z Universe (Maiolino etĀ al., 2023a; Scholtz etĀ al., 2023a; Harikane etĀ al., 2023; Kocevski etĀ al., 2023; Ćbler etĀ al., 2023; Groves etĀ al., 2006). Indeed, at high-z, the environments are systematically more metal poor and the stellar populations are generally younger, increasing the ionization parameter () of star-forming galaxies (SFG). The latter effect, on the one hand, makes SFGs move towards the AGN locus on the BPT; on the other hand the lower metallicity of the Narrow Line Region (NLR) of AGN makes its line ratios move towards and overlap with the star formation locus on these traditional diagnostic diagrams. The detection of high-ionization emission lines (like NāVā, [NeāIV], [NeāV]) can still be considered as a safe tracer for the presence of an AGN (Calabro etĀ al., 2024; Cleri etĀ al., 2023; Brinchmann, 2023; Chisholm etĀ al., 2024), but their detection is still difficult even in the deepest JWST spectroscopic surveys like the JWST Advanced Deep Extragalactic Survey (JADES, Bunker etĀ al., 2023; Scholtz etĀ al., 2023a).
To have a clear picture of the AGN properties and demographics at high-z, it is therefore crucial to identify new AGN selection techniques. Thanks to the large wavelength coverage (m) and sensitivity of JWST it is now possible to have access to a plethora of rest-UV and optical emission lines in high-z galaxies that were not accessible before. Among these lines, multiple JWST spectra have revealed the detection of the [OIII]line. In the metal poor and highly ionized environments of many high-z galaxies, [OIII]is the strongest auroral line, i.e. a collisionally-excited emission line generated from higher energy levels compared to the typical nebular lines. Due to its faintness and proximity to the (generally) stronger Hline, it has been detected and studied almost only in galaxies at , before the launch of JWST. Given its sensitivity to the gas temperature, this line has been extensively used in the local Universe to determine the metallicity of the interstellar medium (ISM) via the so-called ādirect temperature methodā, which exploits the possibility of computing the metallicity by fixing the gas electron temperature derived from the ratio between the [OIII]and the [OIII]nebular lines (see e.g.Ā review by Maiolino & Mannucci, 2019).
However, Brinchmann (2023) first noticed that the anomalously high [OIII]emission in some of the early NIRSpec spectra at z8 could be indicating the presence of an AGN. More recently, Ćbler etĀ al. (2024) have suggested that a strong ratio between this line and the Hline could be a possible indicator for the presence of an AGN. In their work, by using JWST/NIRSpec Integral Field Spectroscopy (IFS) observations, they noticed a spatial correspondence between the AGN broad Hemission and the [OIII]line emission. They also mention the possibility that the enhanced [OIII]emission in correspondence of the AGN could be related to higher ISM temperatures driven by the AGN activity.
An enhancement of the [OIII]emission in the presence of an AGN is qualitatively expected. Indeed, AGN ionizing photons are more energetic than ionizing photons produced in star forming regions on average, hence they can heat the gas more effectively by depositing a larger amount of energy per unit photon. In this work, we will explore this scenario more quantitatively, with the ultimate goal of exploring the possibility of selecting AGN using the [OIII]line both through an observational, empirical approach and via photoionisation models. As a result, we propose three new AGN diagnostic diagrams that can also be effective at high-z.
In SectionĀ 2 we present the local and high-z observational samples and the set of photoionization models considered to test the validity of the new diagnostic diagrams. In SectionĀ 3, we describe the new diagnostic diagrams based on the line ratio [OIII]/Hand the demarcation lines that can be used to separate the AGN population from SFGs. Lastly, in Section 4, we provide interpretations on why these diagnostics work and we show indications on the AGN nature of some newly identified AGN using these diagnostics. Then, we also discuss the implications of this work on the metallicity estimates based on the [OIII]line and on the strong-line metallicity diagnostics.
2 Methods
We selected several AGN and SFG samples and two different sets of photoionization models to populate the diagnostic diagrams that we propose to select AGN. All the observational samples described in the following Sec.Ā 2.1 and Sec.Ā 2.2 are listed in Table 1.
2.1 Low redshift samples
The sample of low-z SFGs is constituted by both normal SFGs and also by the so-called local analogues of high-z galaxies, i.e. local galaxies that show intrinsic properties similar to those observed in high-z galaxies, in terms of metallicities and emission lines ratios (e.g. Izotov etĀ al., 2018, 2019, 2021). The Sloan Digital Sky Survey (SDSS) DR7 (Abazajian etĀ al., 2009) gives the most numerous sample, mostly made of galaxies. Starting from the whole catalog provided by MPA/JHU, we considered only sources with a in all the lines involved in the diagnostic diagrams we are going to present: [OIII], [OIII], [NeāIII], [OII], and H. Then we distinguished between AGN and SFGs using the BPT diagram (Baldwin etĀ al., 1981), taking the AGN demarcation line provided by Kewley etĀ al. (2001) and after considering a further cut in S/N also in the H, Hand [NāII]Ā lines. Furthermore, we exclude from the selection AGN classified in the SDSS catalog as BLAGN, since the catalog does not provide separately the NL and BL components of the Balmer lines and our diagnostics rely on the NL emission only. We also correct all the lines for dust attenuation using the Calzetti etĀ al. (2000) attenuation law. The final SDSS SFG (AGN) sample considered in this work contains () sources.
We also identify several samples of local analogues of high-z galaxies with reported [OIII]line fluxes. From Yang etĀ al. (2017b) we took the sample of so-called āblueberriesā, which are 40 dwarf starburst galaxies with small sizes (kpc), very high ionization ([OIII]/[OII]) and low metallicities (). They were selected from the SDSS DR12 at and followed up spectroscopically using MMT. We also consider 43 āgreen peaā galaxies from Yang etĀ al. (2017a), i.e. nearby SFG with strong [OIII]emission line, 2/3 of them also show a strong Ly emission.
We include in the local analogue sample also a compilation of local SFGs with low metallicities from different works: Izotov etĀ al. (2006), Berg etĀ al. (2012), Izotov etĀ al. (2018), Izotov etĀ al. (2019), Pustilnik etĀ al. (2020), Pustilnik etĀ al. (2021), Nakajima etĀ al. (2022). All these samples were selected considering local sources followed up with different telescopes to be sensitive to the faint [OIII]line, which allowed them to measure the galaxiesā metallicity via the so-called direct method. The metallicities of these sources span the range .
We also include the sample of 165 Extreme Emission Line Galaxies (EELG) reported in AmorĆn etĀ al. (2015), selected from the zCOSMOS survey (Lilly etĀ al., 2007) in the redshift range . These are very compact ( kpc), low-mass () galaxies characterized by specific star formation rates (sSFR) above the main sequence for star-forming galaxies of the same stellar mass and redshift.
The local AGN sample comprises TypeII AGN showing different features and selected for different aims. The most numerous AGN sample is represented by the SDSS sample mentioned above. We then include the Dors etĀ al. (2020) TypeII AGN sample, selected from SDSS DR7 using multiple AGN diagnostic diagrams to guarantee a high purity and showing a range of metallicities between .
The compilation of TypeII AGN includes three sources presented in Seyfert (1943), i.e. NGC1068, NGC1275, and the core of NGC4151. These spectra were obtained at the Mount Wilson Observatory and are among the first galaxies hosting active SMBH ever observed. These three sources clearly show the [OIII]emission line in their spectra.
We consider the sources described in Perna etĀ al. (2017), selected from SDSS spectra associated with X-ray emission and showing outflows, which can be spectrally decomposed from the narrow and broad line region (BLR) emission. We also added both TypeI and TypeII AGN from the S7 survey (Dopita etĀ al., 2015; Thomas etĀ al., 2017), where 131 nearby AGN were primarily selected because of their radio-detection and then followed-up using the WiFeS instrument at high resolution (), allowing to resolve the faint [OIII]line. In Thomas etĀ al. (2017) they provide the narrow components of of the emission lines, already subtracting the broad component, if present.
The 35 AGN selected in Armah etĀ al. (2021) are TypeII AGN in the local Universe () with quite high metallicities (the average metallicity is ), that the authors select to study the neon to oxygen abundance and its evolution with metallicity.
The last sample of local AGN is taken from the MOSFIRE Deep Evolution Field (MOSDEF) survey (Kriek etĀ al., 2015), which targeted the rest-frame optical spectra of H-band-selected Galaxies at . We took the full MOSDEF catalogue and selected only those sources with a [OIII]line with a , and then we distinguished between AGN and SFG using the BPT diagram. With these S/N cuts, we could identify only 3 sources, all classified as AGN. This indicates the difficulty in detecting such a faint line further than in the local Universe in the pre-JWST era.
For all the samples reported above we took dust-corrected fluxes from the works reported in Tab.Ā 1.
2.2 High redshift samples
The High-z samples of AGN and SFGs only comprise sources observed spectroscopically using JWST, the only instrument with enough resolution, sensitivity and IR-coverage to detect and disentangle the [OIII]line at . We considered the full publicly released sample of sources with a medium resolution (MR, ) spectrum in the CEERS (PID: 1345; Finkelstein etĀ al., 2022) and JADES (PID: 1210; Bunker etĀ al., 2023) surveys. For JADES, we considered the emission line fluxes and the AGN selection described in Scholtz etĀ al. (2023a). Starting from a sample of 110 sources with medium resolution (MR) spectroscopy, a reliable redshift and sufficient wavelength coverage, the authors performed a NLAGN selection using multiple AGN rest-frame UV and optical diagnostic diagrams, finally selecting 28 reliable AGN candidates in the redshift range . The diagnostic diagrams involved in the selection include the BPT (Baldwin etĀ al., 1981), the VO87 (Veilleux & Osterbrock, 1987), the CIV/CIII] vs CIII] / HeII and the detection of high-ionization emission lines such as [NeV] [NeIV], NV. To select NLAGN the authors consider in these diagnostics both demarcation lines already defined in the literature, while in other cases (such as the BPT and VOT) they defined new demarcation lines derived considering the distribution of photoionization models (Gutkin etĀ al., 2016; Feltre etĀ al., 2016; Nakajima & Maiolino, 2022). Of all the initial JADES sources, only 19 show a [OIII]detection, 7 of which AGN. We followed the same approach for the CEERS program: starting from 313 MR spectra published by the CEERS collaboration, we selected 217 sources with a reliable spectroscopic redshift and we finally selected 45 AGN candidates (Mazzolari et al., in prep). For this work, we considered 32 sources with the [OIII]line detected among the CEERS sample, 4 of which already classified as AGN. From the same programs, we also collect the BLAGN identified by Maiolino etĀ al. (2023a) and Harikane etĀ al. (2023) at , carefully checking, and in case subtracting, the broad component of the Hline.
We further consider the three galaxies in the galaxy cluster SMACS J0723.3-7327, observed in the Early Release Observations (ERO) JWST programme (Pontoppidan etĀ al., 2022) and analysed in Curti etĀ al. (2023a). Exploiting the [OIII]line detections in the NIRSpec spectra, the authors measured metallicities ranging from extremely metal-poor () to about one-third solar. Two of them were later identified as AGN in Brinchmann (2023) due to the presence of the high-ionization [NeāIV]Ā emission line in one and based on the high ionization parameter () of the other.
The sample of high-z sources presented in Nakajima etĀ al. (2023) includes galaxies from three different early JWST programs (CEERS, ERO, GLASS Treu etĀ al., 2022) at . We selected only the 10 sources with the [OIII]line detected, and we marked as BLAGN those that were selected based on their broad Hline in Harikane etĀ al. (2023). For the latter, we subtract the broad Hcomponent, if present.
We also include in the high-z AGN sample the BLAGN reported in Ćbler etĀ al. (2023) at , whose JWST/NIRSpec Integral Field Spectrograph (IFS) observation shows very high ionization lines and low metallicity (1/4 solar). We also include both the components of the dual AGN candidate at reported in Ćbler etĀ al. (2024). For this source, thanks to IFS spectroscopy, the authors were able to study the displacement between the position of the Hbroad line region (BLR) and the strong [OIII]emission line centroid, which were interpreted as the emissions coming from two distinct sources 620pc apart from each other. Interestingly, the authors also found an almost perfect alignment between the [OIII]emission line peak and the peak of the HBLR, suggesting a correlation between the intensity of the [OIII]line and the presence of an AGN.
We also consider the BLAGN reported in Kokorev et al. (2023) at from JWST UNCOVER Treasury survey (Labbe et al., 2021), showing a robust Hbroad component and an unprecedented black hole to host galaxy mass of at least . An even larger black hole to host galaxy mass ratio was more recently found in the BLAGN analyzed in Juodžbalis et al. (2024b), representing an extreme example of dormant SMBH at high-z. We also considered the BLAGN, triply imaged and lensed by the galaxy cluster Abell2744-QSO1, reported by Furtak et al. (2023). This source is probably undergoing a phase of rapid SMBH growth and is also heavily obscured, since the authors measured an from the Balmer decrement.
Note that in all these TypeI AGN cases (actually most of them type 1.5-1.9) the broad component of the Balmer lines is not detected in H, hence the narrow component of H can be easily measured.
We added to our high-z sample also the extreme SFG reported in Topping etĀ al. (2024). Using distict UV transitions the authors found an electron density , a metal poor ionized gas (i.e. ), and a based on the ratio [OIII]/[OII]. The authors ultimately classified this source as a SFG, based on the ratios between the rest-UV emission lines, but they also found some indications of possible AGN activity, like broad Hcomponent accounting for the of the line flux.
We finally included in the sample also GN-z11, an exceptionally luminous galaxy at z = 10.6. Its JWST spectrum reveals the presence of an AGN through the detection of semi-forbidden lines tracing very high densities (inconsistent with the ISM but typical of the Broad Line Region), other transitions typical of AGN, fast outflows, ionization cones, and a larger Ly halo, consistent with those seen in lower redshift quasars (Maiolino etĀ al., 2024b, 2023b; Scholtz etĀ al., 2023b).
| Sample/References | Type | N included | Redshift | |
| SDSS DR7 Abazajian etĀ al. (2009) | SFG & AGN | 3188 (2377/811) | ||
| Yang etĀ al. (2017b) | Blueberries Galaxies | 40 | ||
| Yang etĀ al. (2017a) | Green Pea Galaxies | 43 | ||
| Low-OH compilation | Low-metallicities SFG | 489 | ||
| AmorĆn etĀ al. (2015) | Extreme Emission Line Galaxies | 165 | ||
| Seyfert (1943) | AGN | 3 | ||
| Dors etĀ al. (2020) | AGN | 26 | ||
| Perna etĀ al. (2017) | AGN | 44 | ||
| S7 Survey Dopita etĀ al. (2015) | AGN | 313 | ||
| Armah etĀ al. (2021) | AGN | 35 | ||
| Kriek etĀ al. (2015) | AGN | 3 | ||
| CEERS (R1000) Finkelstein etĀ al. (2022) | SFG & AGN | 32 (28/4) | ||
| JADES (R1000) Bunker etĀ al. (2023) | SFG & AGN | 19 (12/7) | ||
| Curti etĀ al. (2023b) | SFG & AGN | 3 | ||
| Nakajima etĀ al. (2023) | SFG & AGN | 10 | ||
| Topping etĀ al. (2024) | SFG | 1 | ||
| Ćbler etĀ al. (2023) | AGN | 1 | ||
| Ćbler etĀ al. (2024) | AGN | 2 | ||
| Juodžbalis et al. (2024b) | AGN | 1 | ||
| Kokorev etĀ al. (2023) | AGN | 1 | ||
| Furtak etĀ al. (2023) | AGN | 1 | ||
| Maiolino etĀ al. (2024b) | AGN | 1 |
2.3 Photoionization models
The observational data are compared with the results coming from two different sets of photoionization models. The first set of models was initially described in Gutkin etĀ al. (2016) (for star formation), and Feltre etĀ al. (2016) (for narrow-line AGN emission) and updated with more recent stellar spectra and with a better description of AGN cloud microturbulence (Mignoli etĀ al., 2019; Hirschmann etĀ al., 2019; Vidal-GarcĆa etĀ al., 2024). The photoionisation models are built using the Cloudy code (version c13.03) (Ferland etĀ al., 2013) for star formation and AGN narrow-line regions, assuming a wide range of parameters. In particular they consider gas metallicities (, where ) in the range , ionization parameters of the ionizing source , dust-to-metal ratios (), hydrogen gas densities (), different carbon over oxygen abundances (C/O) and two different initial mass functions (IMF) for star formation models. For a complete list of the values of the different parameters we refer to Table 1 of Feltre etĀ al. (2016). These models include predictions for both the AGN and SFG populations; the main feature differentiating between the AGN and SFG models is the ionizing spectrum, i.e. the spectral energy distribution (SED) of the incident ionizing radiation, the first showing a harder radiation field.
As reported above, the photoionisation models of Gutkin etĀ al. (2016) and Feltre etĀ al. (2016) cover a wide range of parameters and might actually include physical conditions rarely found in the general populations of galaxies and in the local analogs. For example, previous theoretical and observational studies have suggested a correlation between the metallicity and ionization parameter in SF regions (e.g., Dopita etĀ al., 2006; Mingozzi etĀ al., 2020; Ji & Yan, 2022), which implies certain combinations of metallicities and ionization parameters should be rarely, if not never, observed in SF regions. Since the demarcation lines in the new diagrams we propose rely on both the observational data and photoionisation models, we performed the following check to make sure that the part of the models with highly unphysical combinations of parameters does not impact our results.
We performed a likelihood analysis comparing the observed line ratios in our samples of SFG and those predicted by the SFG of Gutkin etĀ al. (2016), in order to limit the inclusion of highly improbable or unphysical combinations of the parameters. This method has been already adopted in making model-based inferences for local galaxies (e.g., Blanc etĀ al., 2015; Mingozzi etĀ al., 2020; Ji & Yan, 2022).
We considered a total of three sets of line ratios, including [OIII]/H, [OIII]/H, and [OII]/H. The inclusion of collisionally excited lines from different ionization states of oxygen as well as recombination lines of hydrogen helps break the degeneracy between the metallicity and ionization parameter in the models. Following the formalism of Blanc etĀ al. (2015), we calculated the likelihood of each model given each data point, and combined the likelihood with a flat prior in the log space spanned by the metallicity and ionization parameter to obtain the posterior distribution of these two parameters. From the posteriors, we obtained the weighted average metallicities and ionization parameters for all SFGs. We then selected a region in the metallicity-ionization parameter space not populated by the posteriors of our sample of SFGs. We note that while the models of Gutkin etĀ al. (2016) and Feltre etĀ al. (2016) have a large set of parameters, the primary drivers of the variations in the predicted line ratios are the metallicity and the ionization parameter. To verify this point, we repeated the above calculation for models with different IMFs, densities, dust-to-gas ratios, and carbon-to-oxygen abundance ratios, which generally resulted in a similar distribution in the metallicity-ionization parameter space. With these results, we adopted a conservative cut to select realistic SFG models based on the percentile of the inferred distribution in the metallicity-ionization parameter space that produces the highest boundary111The boundary is determined by the common outer envelopes of SDSS SFGs and local analogs, as these two samples differ significantly in number and occupy different regions in the parameter space..
Our final cut removes the region of the SFG parameter space spanned by and according to the following relations:
and
In Fig.Ā 5, we plot our fiducial cut together with the inferred distribution of the metallicities and ionization parameters for our selected SFGs based on different sets of models. The cut basically removes regions having very high ionization parameters and high metallicities at the same time. From a physical point of view, due to the dependence of the ionizing spectra of young stellar populations and the mechanical feedback from young stellar populations on the metallicity, it is difficult to maintain a very high ionization parameter at a high enough metallicity (Dopita etĀ al., 2006; Carton etĀ al., 2017). While these physically motivated cuts do not impact the location of the AGN photoionization models in the diagnostic diagrams, the parameter space occupied by SFG models is partially reduced. We emphasize that 99.9% of the considered SFGs lie outside of the excluded parameter space, even assuming different values of , and .
We also compare the emission line ratios of the observational samples with the models described in Nakajima & Maiolino (2022). Using Cloudy, the authors investigated the emission line ratios of SFG, AGN, Population III stars, and Direct Collapse black holes (DCBHs) using the BPASS stellar population models (Eldridge etĀ al., 2017) and a range of other parameters (similar to those considered in Feltre etĀ al. (2016)) listed in Table 1 of Nakajima & Maiolino (2022). In this work, we only consider the models for SFG and AGN, as the very low metallicity values of the other two models () represent extreme scenarios that are not representative of the sources we want to identify with this analysis. In particular, for these models, the authors considered values of and values of from -0.5 to -3.5. Since SFG models of Nakajima & Maiolino (2022) always cover regions of the diagnostic diagrams already covered by the physically-motivated limited Gutkin etĀ al. (2016) and Feltre etĀ al. (2016) SFG models, confirming the goodness of our parameters cuts, we did not apply any cut in this case.
3 Results
In this section, we present the three new diagnostic diagrams proposed to select AGN using the [OIII]line. In these diagrams we combine the auroral line, sensitive to temperature, with various other lines which are sensitive to other additional properties of the ISM, such as ionization parameter and shape of the ionizing SED, hence can further help to disentangle the source of excitation. Additionally, these lines are accessible with JWST at high redshift, and most of them are also close in wavelength, so little affected by dust reddening.
3.1 [OIII]Hvs [OIII][OII]
The diagnostic presented in Fig.Ā 1 is based on the ratio between the [OIII]/Hcompared to [OIII]/[OII]. Hereafter [OII]refers to the sum of [OII]. In the left panel, where we plot the observational samples described in Sec.Ā 2, it is noticeable that the distribution of AGN and SFGs from SDSS are well separated, with AGN occupying a region with higher [OIII]/Hratios compared to normal SFGs. The distribution of the local analogues in this diagram seems to extend the distribution of the local SDSS galaxies towards higher [OIII]/Hand higher [OIII]/[OII], along the diagonal of the plot. On the contrary, the local AGN samples cover a wider area, mostly above the distribution of local analogs and SFGs.
High-z sources are generally distributed in the upper part of the diagnostic with respect to the local samples. We note that a non-negligible fraction of high-z sources not classified as AGN lies in the region of the diagnostic populated by local AGN, while there are also different high-z AGN falling in the same region covered by local analogues.
On the right panel, we show the distribution of the photoionization models presented in Sec. 2.3 according to the same line ratios. We see here that SFG models are predicted to occupy a limited region of this diagram, whose upper boundary almost perfectly corresponds with the distribution of SDSS SFG and local analogs, while AGN are expected to cover both the region occupied by SFGs and also the part of the diagnostic covered by the local AGN samples. Therefore, SFG models are not able to reach the part of the diagram occupied by local AGN, as expected also from the distribution of the observational samples. The opposite is not true, since AGN can occupy the region populated by SFGs, as observed for some local and high-z AGN. These considerations remain true even considering the full grid of SFG models computed in Feltre etĀ al. (2016), as shown in Fig. Ā 7.
As we will discuss, to a first approximation, the ratio [OIII]/[OII]traces the ionization parameter. The fact that, at a given [OIII]/[OII], AGN photoionization can reach much higher values of [OIII]/Hthan photoionization from hot stars is likely a consequence, on average, of the much higher energy of ionizing photons produced by AGN; therefore, at a given ionization parameter, AGN can heat the gas much more effectively than hot stars.
It is worth noting the position of some high-z sources in this diagnostic. The AGN reported in Kokorev etĀ al. (2023) shows an extremely strong [OIII]line and it is placed in a region that is not covered by any model, which probably suggests a very high electron temperature and , as also pointed out in Kokorev etĀ al. (2023). Another significant high-z AGN falling in the AGN-dominated region of this diagnostic is the source corresponding to the BLR centroid in the dual source described in Ćbler etĀ al. (2024), where, in the IFS map, the peak of the broad-line emission spatially corresponds to the peak of the [OIII]line emission. The other component, the [OIII]centroid, showing a fainter [OIII]line emission, is close to the border between the local analogues and the local AGN distribution, together with most of the high-z BLAGN. The AGN region is also populated by different AGN selected among the CEERS and JADES MR NIRSPEC spectra, but we also have different sources not selected as AGN but falling in the AGN-dominated region, as pointed out above. We demonstrate in Sec. 4 that, from the stack of their spectra, we have an indication that some of these sources may host a BLAGN too.
An interesting source in this diagnostic is the SFG reported in Topping etĀ al. (2024). The source shows both strong [OIII]/Hand [OIII]/[OII]emission line ratios, following the extrapolation of the distribution of the local analogues, but at very high values of the ionization parameters and at low metallicities (see Fig.Ā 6), as it is indeed found by the analysis performed in Topping etĀ al. (2024). Looking at the photoionization models, that region is not covered by the models of Gutkin etĀ al. (2016) and Feltre etĀ al. (2016) (not even taking the whole parameters grid), but is instead covered by some AGN models according to Nakajima & Maiolino (2022), potentially suggesting that the main source of ionising radiation in this extreme galaxy is ambiguous.
It is worth noting that the effect of dust reddening on the line ratio [OIII]/[OII], whose components can suffer a different amount of attenuation due to their distance in wavelengths, would move data points towards the right. This can potentially shift AGNs into the SFG+AGN region under significant reddening, but not the opposite, i.e. galaxies dominated by star formation would not contaminate the AGN-only region.
3.2 [OIII]Hvs [NeāIII][OII]
The diagnostic presented in Fig. 2 is based on the ratio between the [OIII]/Hcompared to [NeāIII]/[OII]. Not surprisingly, the distribution of the local and high-z sources in this diagnostic is very similar to the one described in the previous one. Indeed, [NeāIII]/[OII]correlates closely with [OIII]/[OII]both in AGN and SFGs, as shown by Witstok etĀ al. (2021), and as expected given that Neon and Oxygen are both -elements and given that Ne+ and O+ have similar ionisation potentials. The disadvantage of [NeāIII]relative to [OIII]is that it is typically much fainter. The advantage of [NeāIII]is that it is at a shorter wavelength, hence observable at higher redshift (out to z12.3 with NIRSpec), and is closer in wavelength to [OII], hence the [NeāIII]/[OII]ratio is much less affected by dust reddening. As for the previous diagram, normal SFGs and local analogs are distributed in the lower-right part of the diagram, in the same region covered by SFG models, while AGN can be distributed over a wider area, including the upper-left part that cannot be reached by any SFG models or samples. As for the diagnostic described in Sec.Ā 3.1, the separation between the two populations holds even considering the whole grid of the parameter space of the Gutkin etĀ al. (2016) star forming models, as presented in Fig.Ā 6.
Even if the local analogues are less numerous than in Fig.Ā 1 (simply because there are fewer reported [NeāIII]fluxes in the literature), their distributions seem to be again an extrapolation of the distribution of local SFGs towards higher and lower metallicities (see Fig.Ā 6).
The sources that were clearly in the AGN region in the previous diagnostic are still above the SFG distribution in this diagram, and the sources not identified as AGN but lying in the AGN region of the diagnostic in Fig.Ā 1 are in the same region also in Fig.Ā 2. In addition to this latter sample, in Fig.Ā 2 there are few other sources from the CEERS and JADES samples that were not in the diagnostic of Fig.Ā 1, because the [OIII]line was not available, due to the presence of a detector gap. The AGN region is also populated by three additional sources reported in Nakajima etĀ al. (2023) and not classified as AGN in the literature. Their location, however, is still close to the border of the SFG distribution.
The diagnostic diagram presented in this section has some similarity with the so-called āOHNOā diagram that involves the [OIII]/Hvs [NeāIII]/[OII]line ratios (Backhaus etĀ al., 2022, 2023; Zeimann etĀ al., 2015; Cleri etĀ al., 2023; Trump etĀ al., 2023; Feuillet etĀ al., 2024; Killi etĀ al., 2023). The latter is mainly an ionization-sensitive diagram (ionization parameter, shape of the ionizing continuum), but it is also sensitive to metallicity and indeed at high redshift can suffer from low-metallicity galaxy contamination (Scholtz etĀ al., 2023a). The O3Hg vs Ne3O2 diagnostic is instead more stable, also at high redshift, as the separation between the AGN and SFG populations is principally based on a different gas temperature, as we will discuss further in Sec.Ā 4.
3.3 [OIII]Hvs [OIII][OIII]
The third diagnostic is presented in Fig.Ā 3 and is based on the ratio between the [OIII]/Hcompared to [OIII]/[OIII]. This diagnostic was reported also in Ćbler etĀ al. (2024). Looking at the distribution of SDSS SFGs and AGN in the left panel, we note that there is more overlap between the two populations with respect to the previous diagnostics, but there is still a region, characterized by high values of [OIII]/Hand high [OIII]/[OIII], that is populated only by local AGN. The local analogs distribution seems again to be an extrapolation of the trend of the SDSS SFG distribution. In this diagram, there are very few high-z galaxies that are in the AGN-only region of the diagnostic, while most of them are shifted towards lower values of [OIII]/[OIII], even in a region that is not covered by the current photoionization models, as it is evident from the right panel. This trend is emblematically represented by the two extreme AGN reported in Kokorev etĀ al. (2023) and in Furtak etĀ al. (2023). Given the strength of their [OIII]line emission, they are placed in the upper-left part of the plot in a region that cannot be reproduced by any model. The same trend is shown by most of the galaxies that were not selected as AGN, but that lie in the AGN region of the two diagnostics presented above. The distribution of the SFG models matches the distribution of the local galaxies, in particular in its boundary with the AGN-only region, while AGN models cover both the region occupied by SFGs and the one populated only by AGN samples.
In this diagram, assuming the whole parametersā grid of the Feltre etĀ al. (2016) models, we would have a complete overlap between the SFG and AGN models. However, the upper left part of the diagnostic would be covered by SFG models characterized by high values of and high metallicities, a condition not observed in any of our galaxy samples (see AppendixĀ A) and in general in the literature (Kaasinen etĀ al., 2018; Ji & Yan, 2022; Grasha etĀ al., 2022).
3.4 Defining the locus of AGN-only objects
For the three diagnostic diagrams described above, we were able to trace clear demarcation lines providing a separation between the AGN-dominated regions and the regions where SFGs and AGN can overlap. To do so, we considered both the photoionization models and the observational sample distributions. The demarcation lines for the O3Hg vs O32 diagnostic diagram are defined as follows:
Similarly, for the O3Hg vs Ne3O2 diagnostic diagram, the demarcation lines are:
Finally, the demarcation line for the O3Hg vs O33 diagnostic diagram is the following:
These demarcation lines ensure that the fraction of contaminating SDSS SFG in the AGN part of the diagnostics is less than 1-2%, while none of the local analogs lie in the AGN-only part of the three diagnostics.
Such a tiny fraction of contaminants can probably be associated with sources that are not identified as AGN by the BPT diagram (used to distinguish between SDSS SFG and AGN in this work), but that host active SMBHs. The presence of AGN contaminants among BPT-selected SFG and local analogs has already been demonstrated by recent works (Svoboda etĀ al., 2019; Harish etĀ al., 2023).
Furthermore, it is important to note that the AGN selection based on these demarcation lines is not a necessary condition for a source to be selected as an AGN, but rather a sufficient condition: objects above the demarcation line are AGN-dominated with high confidence, while objects below the demarcation line can either be SFGs or AGN.
4 Discussion
By comparing the distribution of the observational samples with the distribution of the photoionization models, we can explain why these diagnostic diagrams can separate part of the AGN population from SFGs. In FigĀ 6, we plot the AGN and SFG models in the same diagnostic diagrams presented in Sec.Ā 3, but highlighting the variation of and , i.e. the two main parameters affecting the distribution of the models in these plots. We know that, since they are very local, SDSS AGN and SFG samples are populated mostly by solar metallicity and moderate-to-low ionization parameter sources, as shown also by the models occupying the same region of these samples in FigĀ 6. On the contrary, local analogs are characterized, on average, by lower metallicities and higher ionization parameters (as high-z galaxies). Indeed, in Fig.Ā 1 and Fig.Ā 2, they fill the regions that extend from the SDSS SFG towards higher [OIII]/Hand higher [OIII]/[OII]or [NeāIII]/ [OII], in close agreement with the trend of photoionization models for higher and in the same diagnostics (see Fig.Ā 6).
The reason for the presence, in particular in Fig.Ā 1 and Fig.Ā 2, of a region that is populated only by AGN, at high [OIII]/Hvalues, can be explained in the following way. Since the [OIII]line is a collisionally excited line generated from high energy levels, its intensity relative to His directly related to the temperature and, secondly, to the metallicity of the ISM and ionization parameter. The main difference between the AGN and SFGs is the SED of the incident ionizing radiation. For AGN the ionizing radiation is given by the emission of the accretion disk, while for SFGs it originates from young star clusters. The former are characterized by a harder spectrum and, therefore, on average photons are more energetic (see Fig.1 in Feltre etĀ al., 2016). Higher energy per ionizing photon, deposited into the ISM, results into a higher effective heating, hence higher electron temperature (hence higher [OIII]) at a given ionization parameter.
4.1 Stacking of the AGN candidates
To further demonstrate the effectiveness of the diagnostic diagrams proposed in this work we explored potential tracers of AGN activity in those high-z galaxies lying in the AGN region of the diagnostics diagrams in Fig. 1 and Fig. 2 but not previously identified as AGN based on the selection performed in Scholtz etĀ al. (2023a) and Mazzolari et al., in prep.
Our stacking procedure is as follows. We first shift the spectra to the rest-frame and normalise them to the peak of the Hline, that is available for all the sources involved. This ensures that the bright sources are not dominating the final stacks. We then resample each of the spectra to the common wavelength grid. We verified that the wavelength grid does not impact our conclusions by repeating the analysis using wavelength bins ranging from the best (km/s) to the worst (km/s) resolution. Finally, before stacking the spectra, we fit and subtract the continuum from the rest-frame rebinned spectra. We performed an inverse variance stacking of 15 sources (11 from CEERS and 4 from JADES).
We first fit simultaneously the Hand [NāII]Ā wavelength region using only a single Gaussian profile per emission line, with common FWHM and redshift for all three emission lines. We fixed the [NāII]Ā doublet ratio to be 3. We show the narrow emission line fit in the top left panel of Fig. 4. We detect broad residuals around the Hemission line. These broad components can arise from either a broad line region or an outflow.
To model the broad wing in the H, we refit the Hand [NāII]Ā with two additional models - BLR model (by fitting one additional broad Gaussian component to Honly) and an outflow model by adding broad Gaussian components to both Hand [NāII]Ā with fixed FWHM and centroid. Using the Bayesian Information Criterion (BIC) parameter (Liddle, 2007), we found that the BLR model is strongly preferred to the narrow-only fit (). The broad component has an S/N of 10, making this a solid detection. On the contrary, the fit with the outflow model did not return any broad component in the [NāII]Ā lines.
We then used the same approach with the Hand [OāIII]ācomplex, to test the possibility that this broad component could be associated with an outflow on larger scales, determining a broad [OIII]. In the fit shown in the last panel of Fig.Ā 4, we imposed the broad component of the [OāIII]āto have the same kinematics as the broad component of the H, i.e. same FWHM and velocity offset. However, we didnāt find a significant broad line, and therefore outflow, detection in the [OāIII]ā. In particular, by imposing a broad outflow component to the [OāIII]āand comparing with the NL-only fit, it gives a , meaning that the two models are almost equally preferred, hence a broad [OIII]component is not required.
The absence of a counterpart in (brighter) [OIII]strongly disfavours an outflow origin of the broad component of H. Therefore, the detection of a broad component in the Hemission line suggests the presence of an AGN BLR in the spectra of at least some of these galaxies, which is too weak to be detected in the individual spectra. This indicates that galaxies located in the upper region of our diagnostic diagrams, even if originally not classified as AGN, actually may host faint TypeI AGN. This result provides confidence that the diagnostic diagrams presented in this work are effective in identifying AGN above our proposed demarcation lines, although they cannot distinguish between AGN and SFGs below the demarcation line.
4.2 Ruling out [FeII] contamination
We investigate the possibility that the [OIII]emission line in the high-z sources located in the AGN-only region of the diagnostics could be contaminated by the [FeII] emission line (Curti et al., 2017; Arellano-Córdova et al., 2022). Most of the high-z [OIII]detections come from JWST MR spectra, whose resolution is not enough to disentangle the two lines. Curti et al. (2017) showed that this contamination is possible, but most likely in high-metallicity galaxies, while our high-z sample is generally metal poor.
To investigate the possible iron contamination of the [OIII]line, we consider the close [FeII] emission line, which is isolated and originates from the same energy level as the [FeII]. We ran a grid of Cloudy photoionization models over a wide range of metallicities (with ) and ionization parameters (with ), and found that the flux ratio of [FeII] / [FeII]222Our calculation takes into account 6 lines from blended at 4288 Ć
Ā and 8 lines from blended at 4360 Ć
, although the majority of the fluxes is contributed by [FeII] and [FeII]. is roughly a constant of . Therefore, we can use the [FeII] emission line to reliably constrain the intensity of the [FeII]. We did not detect the [FeII] line in any of the high-z sources in the AGN-only regions of the diagnostics. Also performing a weighted spectral stacking of all the spectra in the same region, the [FeII] emission line remains undetected. Therefore we conclude that any possible contribution to the [OIII]detections given by the [FeII] emission line is negligible.
4.3 Impact on metallicity estimates and strong-line metallicity diagnostics
The [OIII]emission line has been widely used in the literature to derive metallicities via the so-called method (see Maiolino & Mannucci, 2019, for a review), although Marconi etĀ al. (2024) have shown that these metallicities may be biased low. The metallicities inferred from the method have then been used to calibrate the so-called strong-line metallicity diagnostics, i.e. ratios of (optical) emission lines typically much stronger than the auroral lines, which can be used to estimate the metallicities on larger samples of galaxies. These calibrations have been inferred both by using local samples of galaxies (e.g. Maiolino etĀ al., 2008; Bian etĀ al., 2018; Curti etĀ al., 2020) and, especially with JWST, at high redshift (Sanders etĀ al., 2023; Laseter etĀ al., 2024).
In most of these past studies care was taken in using only SFGs, or assuming that the gas ionization was dominated by young hot stars. The finding of consistent populations of AGN at high redshift by recent JWST studies, and especially our own finding in this paper that a significant number of strong [OIII]detections at high redshift are due to AGN heating, may appear as concerning when using this transition to infer the gas metallicity. However, the [OIII]line (together with [OIII], or [NeāIII]as a proxy when the latter is not available) provides a measurement of the temperature in the O+2 zone, regardless of the nature of the ionizing source. So its reliability remains unaffected in the case of photoionization by AGN (modulo the potential biases discussed in Marconi etĀ al., 2024, which anyway are a concern also for SFGs). What is obviously important is to properly use this information to infer the metallicity, and in particular by applying the proper ionization correction factors, which are different in the case of AGN and SFGs. In particular, once the abundance of is derived, one has to estimate the contribution to the abundances from and . The former is very difficult to assess, even in AGN, as there are no [OIV] strong transitions in the wavelength ranges typically accessible. Yet, Dors etĀ al. (2020) estimate that, even in AGN, the contribution of is typically negligible. The contribution from is estimated from the [OII]doublet; however, the key issue is that the temperature of the region is different from the temperature of the region probed by the [OIII]line. If [OII] auroral lines are not available (as in the vast majority of the cases at high redshift), then one has to assume a relation between and . This relation has been extensively explored in local SFGs, however it cannot be applied to AGN. Indeed, assuming in AGN the same temperature relation as for SFGs results into a large underestimation of the metallicities (Dors etĀ al., 2015). This issue is greatly mitigated once AGN-bespoke relations between and are adopted (Dors etĀ al., 2020).
In summary, the fact that at high-redshift [OIII]is boosted by AGN heating does not prevent it from being used for measuring the metallicity in those galaxies, provided that the proper ionization correction factors are adopted and, in particular, the adequate and is adopted (in absence of [OII] auroral lines).
However, the presence of AGN ionization, excitation and heating, certainly affects the empirical calibrations of the strong line diagnostics. Indeed, at a given gas temperature, the AGN radiation (with different ionizing shape and, typically, higher ionization parameter) results in a different ionization structure of the gas clouds, and also different collisional excitation rates of the various transitions typically used in the strong line diagnostics ([OIII], [OII], [NeIII], [NII], [SII]). As a consequence, separate and different empirical metallicity calibrations should be inferred for SFGs and AGN host galaxies. This has been attempted in the local universe (Dors, 2021). However, the same effort should be undertaken at high redshift, given the large abundance of AGN. Inferring empirical strong-line metallicity calibrations differentiating AGN and SFGs will be the focus of a separate dedicated paper.
5 Conclusions
While JWST has revealed that some of the classical AGN diagnostics break down at high redshift, it has also opened the opportunity to explore new diagnostics. In this work, we studied the possibility of selecting AGN using the [OIII]auroral line, whose detection in large numbers of high-z galaxies has become possible with JWST. In particular, we proposed three new diagnostic diagrams, and three corresponding demarcation lines, that allow to identify a large population of AGN from SFGs, by providing a sufficient (but not necessary) condition to claim the presence of an AGN.
To demonstrate the effectiveness of these diagnostics, we used multiple observational samples of local and high-z SFGs and AGN as well as photoionization models from Feltre etĀ al. (2016) and Nakajima & Maiolino (2022).
We specifically proposed the following three diagnostic diagrams:
ā¢
[OIII]/H vs [OIII]5007/[OII]3727. This is the most thoroughly explored diagram from the empirical perspective (it has the largest observational test sample), and can be used with NIRSpec out to z9.4.
ā¢
[OIII]/H vs [NeIII]3869/[OII]3727. This diagram is the least sensitive to dust extinction given the wavelength proximity of both line ratios. Additionally, it can be used with NIRSpec out to z10.9.
ā¢
[OIII]/H vs [OIII]5007/[OIII]. Among the three, this is the one with the smallest region where AGN can be safely identified, but it is also relatively dust insensitive and requires the detection of only three lines instead of four, making it applicable to larger samples.
In each of these cases, we have provided a demarcation line above which (i.e. with higher [OIII]/H values) objects can be safely identified as hosting an AGN that dominates the nebular emission lines excitation.
We have illustrated that at least some of few objects falling into the AGN-only region and not previously identified as AGN, do show AGN signatures when stacked. This further supports the effectiveness of the diagnostics. At the same time, we did not find any indication of [FeII] contamination of the [OIII]emission line in these sources, which could in principle artificially increase the [OIII]line due to blending effects.
The physics behind these diagnostics is tightly linked to the primary property of AGN, i.e. the hardness of their ionizing spectrum. The average energy of AGNās ionizing photons is much higher than the energy of hot, young stars in SFGs. Therefore, at a given ionizing radiation field intensity, AGN photons are more effective in heating the ionized gas than in SF regions, hence yielding higher temperature, hence higher [OIII]relative to H (which normalizes by the overall photoionizing radiation field).
We stress that being above the demarcation line in the proposed diagrams is a sufficient but not necessary condition for an object to be identified as an AGN. Galaxies located below such demarcation lines can be powered by star formation and/or AGN. At the same time, the contamination from SFG is almost completely negligible in the upper part of the diagnostics, as demonstrated by the distribution of low-redshift observational samples, photoionization models. Moreover also the few high-redshift ambigous cases in the AGN-only region turn out to be AGN based on the stacking.
Finally, we note that the fact that strong auroral lines are often associated with AGN does not imply that they cannot be used for direct metallicity measurements (provided that proper ionization corrections are applied), but it does affect the calibration of strong line metallicity diagnostics, calling for new, AGN-specific, calibrations of these diagrams.
Acknowledgements
We thank the anonymous referee for useful suggestions which improved the quality of the paper.
GM acknowledges useful conversations with Roberto Gilli, Marcella Brusa, Sandro Tacchella, William Baker, Callum Witten, Bartolomeo Trefoloni, Lola Danhaive, Ignas Juodzbalis, Amanda Stoffer, Brian Xing Jiang, and William McClaymont.
FDE, JS and RM acknowledge support by the Science and Technology Facilities Council (STFC), from the ERC Advanced Grant 695671 āQUENCHā, and by the
UKRI Frontier Research grant RISEandFALL. RM also acknowledges funding from a research professorship from the Royal Society.
HĆ gratefully acknowledges support by the Isaac Newton Trust and by the Kavli Foundation through a Newton-Kavli Junior Fellowship.
Appendix A Cut in photoionization models
In Fig.Ā 5 we show the distribution of the SDSS SFG and local analogues samples in the vs plane as derived following the likelihood procedure with the Gutkin etĀ al. (2016) models described in Sec.Ā 2.3. Even considering different values of the C/O, , and parameters in the models, the chosen excluded region contains less than % of the sources (located very close to the borders).
Appendix B Photoionization models properties
In Fig.Ā 6 we show the distribution of the SFG and AGN photoionization models from Gutkin etĀ al. (2016), Feltre etĀ al. (2016), and Nakajima & Maiolino (2022) in the three diagnostic diagrams presented in Sec.Ā 3. In particular, the plots highlight the variation of and of the models across the diagnostics.
Appendix C Diagnostic diagrams including all photoionization models
The demarcation lines presented in Sec.Ā 3.4 were defined considering both the distribution of the observational samples and of photoionization models with the cut described in Sec.Ā 2.3 and shown in Fig.Ā 5. If we consider the entire grid of parameters for the models of Gutkin etĀ al. (2016) and Feltre etĀ al. (2016), we obtain the distributions shown in Fig.Ā 7. In this case, the demarcation lines for the O3Hg-O32 and O3Hg-Ne3O2 diagnostic plots still hold, except for a very small fraction of SFG models populating the AGN-only part of the two diagnostics. On the contrary, in the O3Hg-O33 diagnostic plot, there is an almost complete superposition between the SFG and AGN models. Even if this superposition is determined by highly unlikely combinations of parameters in the models, we cannot completely rule out the possibility that the AGN-only region of the O3Hg-O33 diagnostic diagram could actually be partially populated also by SFG.